OpenStudy (anonymous):

Reduce to simplest form: 32x^3y ------- 24xy^4 Please Help and show all work Please Please

OpenStudy (anonymous):

Someone please Help Me!!!

OpenStudy (lgbasallote):

\[\large \frac{32x^3y}{24xy^4} \implies \frac{32}{24} \times \frac{x^3}{x} \times \frac{y}{y^4}\] does that help?

OpenStudy (anonymous):

Kinda but still not sure how to solve?

OpenStudy (lgbasallote):

which one?

OpenStudy (anonymous):

x^3 -- x

OpenStudy (anonymous):

And Y --- Y^4

OpenStudy (lgbasallote):

\[\huge \frac{x^m}{x^n} = x^{m-n}\] for example... \[\LARGE\frac{x^3}{x^2} \implies x^{3-2} \implies x^1 \implies x\] does that help?

OpenStudy (anonymous):

Still really confused?